OpenStudy (anonymous):

how do you do a problem like... "If a 10.0g ball of iron at 160.0c is dropped into 50.0g of water at 20.0c in an insulated container, what will be the final temp of the water? the specific hear of iron is .444j/g and that of water is 4.18j/g."?

OpenStudy (anonymous):

Well you will need to convert everything into the same specific heat capacity... 4.18/0.444 = 9.41 Then multiply the mass of the iron by 9.41 to give 94.1 (equivalent grams of water) I would then add the two volumes together (94.1 + 50) to give 144.1 then take 20(degrees) * (50/144.1) = 6.94 and 160*(94.1/144.1) = 104.48 then add them together :111.42 I think that is the answer 111.42degress c however i am not too sure... as your teacher, or whatever, this method might be wrong